40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1556     Entry time: Thu May 7 17:59:23 2009
Author: Alberto 
Type: Configuration 
Category:  
Subject: MC WFS 
This afternoon the MC could not get locked.
I first checked the Osems values at the MC mirrors and compared them to the trend of the last few hours. That showed that the alignment of the mirrors had slightly changed. I then brought each mirror back to its old alignment state.
 
That let the LSC loop lock the MC, although the reflected power was still high (1.5V) and the WFS control wouldn't engage.
 
Since earlier during the day I was working on the AS table, it is possible that I inadvertently hit the MC REFL beam splitter misaligning the beam to the MC WFS.
To exclude that there was a problem in the suspensions, before touching the WFS, I checked that the cables at the MC's ends and those going to the ADC in the rack were well pushed in.
 
Then I proceeded in centering the beam on both the WFS, balancing the power over the QPDs.
 
In the end the MC could lock again properly.
 
ELOG V3.1.3-